OpenStudy (anonymous):

Find g of f if f(x) =2x-1 and g(x)=x+2

OpenStudy (anonymous):

@amistre64 @Nurali

OpenStudy (anonymous):

2x - 3 2x + 1 2x - 1 2x + 3

OpenStudy (anonymous):

g of f is g(f function) = g(2x-1) = 1(2x-1) +2 = 2x+1

OpenStudy (amistre64):

g of f, means that g "is composed of" f